Suppose that f(x)=3x2−4.

(A) Find the slope of the line tangent to f(x) at x=−1.
(B) Find the instantaneous rate of change of f(x) at x=−1.
(C) Find the equation of the line tangent to f(x) at x=−1. y=
1
Expert's answer
2020-11-10T19:56:06-0500

(A) Given: "f(x)=" "3x^2-4"

Require to find the slope of the line tangent to "f(x)" at "x=-1"

Find the derivative of the given function "f(x)" = "3x^2 - 4"

Now "f(x)=" "3x^2 - 4" "\Rightarrow f'(x)=3(2x) - 0 = 6x" (Using Power Rule)

Slope of the line tangent to "f(x)" at "x=-1" is given by

"m=f'(-1)=6(-1)=-6"

Therefore, slope of the line tangent to "f(x)" at "x=-1" is "m=-6"

(B)Require to find the instantaneous rate of change of "f(x)" at "x=-1"

Instantaneous rate of change of "f(x)" at "x=-1" is given by "f'(-1)" and

"f'(-1)=-6"

Therefore, instantaneous rate of change = - 1

(C)Require to find the equation of the line tangent to "f(x)" at "x=-1"

From part(A), we have slope of the line tangent to "f(x)" at "x=-1" is "m=-6"

And "x=-1" "\Rightarrow f(-1)=3(-1)^2- 4 = 3-4=-1"

Now let us find the equation of the line tangent to "f(x)" at the point (- 1, - 1) using point-slope form.

Equation of the line tangent to f(x) in point-slope form is

"y-y_{1}=m(x-x_{1})" , where "(x_{1},y_{1})= (-1, -1)" and "m=-6"

Using the above, equation of the line tangent to f(x) at x = - 1 is

"y-(-1)=-6(x-(-1))"

That is, "y+1=-6(x+1)"

That is, "y=-6x-6-1"

Therefore, equation of the line tangent to the given function is

"y=-6x-7"